The weights of packages filled by machine are normally distributed about a mean of 25 ounces, with a standard deviation of one ounce. What is the probability that n packages from the machine will have an average weight of less than 24 ounces if n = 1, 4, 16, 64? (10)


1
Expert's answer
2021-02-24T06:09:48-0500

Mean, "u=25"

Standard deviation, "\sigma=1"


Average weight, "\bar{X}=24"


Case-1 When n=1


"P(z<24)=1-P(24<z<25) =1-\dfrac{\bar{X}-\mu}{\frac{\sigma}{\sqrt{n}}}=1-\dfrac{24-25}{\frac{1}{\sqrt{1}}}=1-(-1)=2"

The value of z is 0.5080.


Case-2 When n=4


"P(z<24)=1-P(24<z<25) =1-\dfrac{\bar{X}-\mu}{\frac{\sigma}{\sqrt{n}}}=1-\dfrac{24-25}{\frac{1}{\sqrt{4}}}=1-(-2)=3"

The value of z is 0.6852


Case -3 When n=16


"P(z<24)=1-P(24<z<25) =1-\dfrac{\bar{X}-\mu}{\frac{\sigma}{\sqrt{n}}}=1-\dfrac{24-25}{\frac{1}{\sqrt{16}}}=1-(-4)=5"

The value of z is 0.8453


Case-4 When n=64,


"P(z<24)=1-P(24<z<25) =1-\dfrac{\bar{X}-\mu}{\frac{\sigma}{\sqrt{n}}}=1-\dfrac{24-25}{\frac{1}{\sqrt{64}}}=1-(-8)=9"

The value of z is-0.9568